# A Comprehensive Social Bias Audit of Contrastive Vision–Language Models

### v4

@@ Abstract
Vision–language models (VLMs) deliver strong zero-shot recognition but frequently inherit social biases from their training data. We systematically disentangle three design factors—model size, training-data scale, and training-data composition—by comparing CLIP and OpenCLIP, two models that share an identical contrastive objective yet differ in encoder width and in the image–text corpora they consume (400 M proprietary pairs versus 2 B LAION pairs). Across balanced face-analysis benchmarks we find that enlarging the encoder reduces gender skew in CLIP but amplifies both gender and racial skew in OpenCLIP; expanding the corpus from 400 M to 2 B pairs doubles OpenCLIP’s gender bias. When we match model and data budgets, the two corpora expose a fairness trade-off: CLIP-style data incurs more gender bias, whereas LAION-style data incurs more racial bias. These results challenge the intuition that “bigger models or datasets are automatically fairer” and instead spotlight corpus composition as a key driver of bias. We release all code and evaluation scripts to enable transparent, reproducible auditing of future VLMs.
[…]
A Comprehensive Social Bias Audit of Contrastive Vision–Language Models
Zahraa Al Sahili1thanks:   Corresponding author.  Ioannis Patras1  Matthew Purver1,2 1Queen Mary University of London, UK 2Institut Jožef Stefan, Slovenia {z.alsahili, i.patras, m.purver}@qmul.ac.uk
@@ 1 Introduction
Existing audits seldom answer this question because they vary only one factor or inspect a single model family, leaving the effects of model capacity, data volume, and data composition entangled. In language-only settings, scaling can either help or hurt fairness depending on objective and metric (Liang et al., 2021), but systematic evidence for VLMs is still limited.
We address this gap with the first controlled study that simultaneously disentangles three axes while holding the contrastive loss fixed: encoder size (ViT-B/32 vs. ViT-L/14), corpus scale (400 M vs. 2 B image–text pairs), and corpus composition (CLIP’s proprietary dataset vs. the LAION-400M/2B public crawl). Bias is measured on 10 k balanced FairFace images and the PATA social-perception set using a denigration probe (crime/animal) and two stereotype probes (communion, agency) (Hausladen et al., 2024). We report both Max Skew across demographic groups and corpus-level harm rates.
Our findings overturn the intuition that “bigger models or datasets are automatically fairer.” Enlarging the encoder reduces gender skew in CLIP but increases both gender and racial skew in OpenCLIP; scaling the corpus from 400 M to 2 B pairs doubles OpenCLIP’s gender skew; and, at equivalent encoder and data sizes, CLIP incurs more race bias while OpenCLIP remains markedly more gender-biased. These patterns highlight training-data composition—rather than loss design—as a dominant driver of social bias in VLMs.
Contributions.
  • We present a controlled experimental framework that disentangles model size, training-data scale, and training-data composition in CLIP-style VLMs while keeping the loss function constant.
  • We release a public audit of denigration and social-perception bias scores for four widely used open-source models, along with code for reproducible evaluation.
  • We provide empirical evidence that naïve scaling can exacerbate bias, underscoring the need for data-centric mitigation strategies in addition to architectural advances.
By clarifying how architectural and data decisions jointly shape bias, we move toward principled guidelines for building fairer open-vocabulary multimodal systems.
@@ 2 Related Work
We situate our study at the intersection of three longstanding research threads: (i) dataset bias in vision–language models (VLMs), (ii) bias-measurement and mitigation frameworks, and (iii) the scaling literature.
@@ 2.2 Bias-Measurement Frameworks and Mitigation
Task-agnostic toolkits now quantify social bias in VLM outputs. MMBias measures stereotype leakage in captioning and VQA (Janghorbani and de Melo, 2023), whereas DebiasCLIP prepends learned visual tokens to attenuate gender stereotypes (Berg et al., 2022). Embedding-space analyses reveal that CLIP encodes the family $\leftrightarrow$ career gender stereotype (Bianchi et al., 2023; Liang et al., 2021). Prompt-level interventions, such as the Debiasing and VisDebiasing prefixes in ModSCAN, reduce but do not eradicate skew (Jiang et al., 2024). Beyond CLIP, BendVLM performs nonlinear, test-time debiasing of VLM embeddings without fine-tuning (Gerych et al., 2024), Association-Free Diffusion mitigates object-people stereotype transfer in text-to-image generation (Zhou et al., 2024), FairCoT leverages chain-of-thought reasoning in multimodal LLMs to iteratively refine prompts and improve fairness in text-to-image generation (Al Sahili et al., 2024), and VHELM supplies a multi-dimensional benchmark covering bias, fairness, toxicity and safety (Lee et al., 2024). Janghorbani and De Melo (2023) introduce a unified framework for stereotypical bias across modalities, while Raj et al. (2024) expose hidden biased associations that evade existing diagnostics.
@@ 2.3 Scaling Effects on Bias
Scaling model parameters or data volume is often viewed as a route to robustness, yet empirical findings remain mixed. In language models, larger GPT-style systems sometimes reduce toxic generation (Hoffmann et al., 2022) but can entrench occupational stereotypes (Liang et al., 2021). Ghate et al. (2025) show that intrinsic bias is largely predictable from pre-training data, not architecture. Liu et al. (2024) offer a probabilistic framework revealing that stereotype volatility increases with model size. For VLMs, Birhane et al. (2023) warned that expanding LAION from 400M to 5B images risks amplifying existing harms, whereas higher image resolution can mitigate certain gender biases . Our controlled study extends this line by showing that size, scale, and loss composition interact non-linearly: the same parameter increase that softens bias in CLIP amplifies it in OpenCLIP, and a five-fold data increase leaves CLIP’s skew almost unchanged while doubling that of OpenCLIP.
In summary, prior work establishes that VLMs inherit social bias, offers diverse metrics and mitigation heuristics, and yields conflicting evidence on the benefits of scaling. We advance the field by isolating the individual and joint contributions of model size, data scale, and loss decomposition—an experimental control that existing audits lack.
@@ 3 Methodology
[…]
@@
Where a cell is missing we mark it explicitly and confine statistical comparisons to the subset of cells that share all three settings.
[…]
@@ 3.2 Evaluation Datasets
PATA.
[…]
Communion and Agency.
[…]
3.4 Bias Metrics
[…]
$s_{G}(g,c)=\frac{1}{|D_{g}|}\sum_{I\in D_{g}}\cos\bigl{(}v(I),e(c)\bigr{)}.$
We quantify disparity with the Max Skew metric. For any two demographic groups $A$ and $B$ with outcome proportions $p_{A}$ and $p_{B}$,
$S_{A,B}=\max\Bigl{(}\bigl{|}\tfrac{p_{A}-p_{B}}{p_{B}}\bigr{|},\,\bigl{|}\tfrac{p_{B}-p_{A}}{p_{A}}\bigr{|}\Bigr{)}.$
[…]
@@ 3.5 Inference Details
[…]
@@ 4 Results
This section traces how encoder size, pre-training scale, and training-data composition alter social bias. All numbers come from Table 1; Figures 1, 2, and 3 visualise the same effects.
Data Model Sz DSz $\max s_{G}^{c}$ $\max s_{G}^{com}$ $\max s_{G}^{ag}$ $\mu\max s_{R}^{c}$ $\mu\max s_{R}^{com}$ $\mu\max s_{R}^{ag}$ C % NH % NC % NA
@@
Table 2: Bias component deltas across model size, data size, and data decomposition. Red = increase in bias (undesirable), Blue = reduction in bias (desirable).
@@
Table 3: Race-related bias deltas across model size, data size, and dataset decomposition. Red = increase in bias; Blue = decrease in bias.

### v5

# Data Matters Most: Auditing Social Bias in Contrastive Vision–Language Models

@@ Abstract
Vision-language models (VLMs) deliver strong zero-shot recognition but frequently inherit social biases from their training data. We systematically disentangle three design factors—model size, training-data scale, and training-data source——by comparing CLIP and OpenCLIP, two models that share an identical contrastive objective yet differ in encoder width and in the image–text corpora on which they are pre-trained(400M proprietary pairs vs. 400M/2B LAION). Across balanced face-analysis benchmarks, enlarging the encoder reduces gender skew in CLIP but amplifies both gender and racial skew in OpenCLIP; increasing the LAION corpus from 400M to 2B further increases OpenCLIP bias. At matched model and data budgets, substituting proprietary data with LAION improves gender fairness while increasing racial skew, underscoring data source as the primary driver of bias patterns.
We also evaluate three post-hoc, test-time debiasing strategies — Bias Prompts, Prompt Array, and SANER. Debiasing reduces but does not eliminate harm, and its effectiveness is source- and size-dependent: Bias Prompts most effectively reduce gender skew in CLIP at smaller model sizes, whereas Prompt Array and SANER more reliably reduce racial skew in OpenCLIP; scaling LAION reconfigures which method is most fair. Taken together, these findings challenge the assumption that bigger models or datasets are automatically fairer and foreground training data source as the key determinant of both bias and mitigation efficacy. We release code and evaluation scripts to enable transparent, reproducible auditing of future VLMs. [^1] URL redacted for double-blind review
@@ 1 Introduction
Existing audits seldom answer this question because they vary only one factor or inspect a single model family, leaving the effects of model capacity, data volume, and data source entangled. In language-only settings, scaling can either help or hurt fairness depending on the objective and metric (Liang et al., 2021), but systematic evidence for VLMs is still limited.
We address this gap with a controlled study that holds the contrastive objective fixed while disentangling three axes: encoder size (ViT-B/32 vs. ViT-L/14), corpus scale (400M vs. 2B image–text pairs), and corpus source (CLIP’s proprietary dataset vs. the LAION-400M/2B public crawl). Bias is measured on 10k balanced FairFace images and the PATA social-perception set using a denigration probe (crime/animal) and two stereotype probes (communion, agency) (Hausladen et al., 2024). We report both Max Skew across demographic groups and corpus-level harm rates.
Our findings overturn the intuition that “bigger models or datasets are automatically fairer.” Enlarging the encoder reduces gender skew in CLIP but increases both gender and racial skew in OpenCLIP; scaling the corpus from 400M to 2B pairs further amplifies OpenCLIP’s bias; and, at matched model and data budgets, swapping the proprietary crawl for LAION trades improved gender fairness for heightened racial skew. These patterns identify training-data source as a first-order driver of bias in contrastive VLMs.
Beyond auditing baselines, we evaluate three post-hoc, test-time debiasing strategies—Bias Prompts, Prompt Array, and SANER. Debiasing reduces but does not eliminate harm, and its effectiveness is source- and size-dependent: Bias Prompts are most effective for reducing gender skew in CLIP at smaller model sizes, whereas Prompt Array and SANER more reliably reduce racial skew in OpenCLIP. These results argue for principled, scale-stable debiasing objectives and corpus-aware processes—validated at the deployment model size and robust across sources and axes.
Accordingly, our contributions are fourfold:
  • A controlled experimental framework that disentangles model size, training-data scale, and training-data source in CLIP-style VLMs while keeping the loss function constant;
  • A public audit of denigration and social-perception bias scores for widely used open-source checkpoints, with both disparity and harm-frequency reporting;
  • A comparative evaluation of three test-time debiasing strategies across six metrics (gender/race $\times$ crime/communion/agency), showing that mitigation efficacy depends on data source and model size;
  • Code and evaluation scripts to enable transparent, reproducible auditing of future VLMs under identical controls.
By clarifying how architectural and data decisions jointly shape bias—and how debiasing interacts with those decisions—we move toward principled guidelines for building fairer open-vocabulary multimodal systems.
@@ 2 Related Work
We situate our study at the intersection of three longstanding research threads: (i) dataset bias in vision-language models (VLMs), (ii) bias-measurement and mitigation frameworks, and (iii) the scaling literature.
@@ 2.2 Bias-Measurement Frameworks and Mitigation
Task-agnostic toolkits now quantify social bias in VLM outputs. MMBias measures stereotype leakage in captioning and VQA (Janghorbani & de Melo, 2023), while DebiasCLIP reduces gender stereotypes by prepending learned visual tokens (Berg et al., 2022b). Embedding-space analyses further show that CLIP encodes the family $\leftrightarrow$ career gender stereotype (Bianchi et al., 2023; Liang et al., 2021). Prompt-level interventions—e.g., the Debiasing and VisDebiasing prefixes in ModSCAN—can attenuate skew but rarely eliminate it (Jiang et al., 2024).
Within test-time textual controls, three complementary approaches have emerged. Biased Prompts removes protected-attribute directions from CLIP text features via a calibrated, closed-form projection, reducing leakage without extra training or labels (Chuang et al., 2023). Prompt Array adversarially learns a small set of tokens prepended to sensitive queries, aiming to suppress bias while preserving image–text alignment (Berg et al., 2022a). SANER provides an annotation-free societal attribute neutralizer that targets only attribute-neutral text features, preserving explicitly specified attributes through neutralization with lightweight modifiers and reconstruction/contrastive regularizers (Hirota et al., 2025). Beyond CLIP, BendVLM performs nonlinear, test-time debiasing without fine-tuning (Gerych et al., 2024); Association-Free Diffusion mitigates object–people stereotype transfer in text-to-image generation (Zhou et al., 2024); FairCoT leverages chain-of-thought prompting in multimodal LLMs (Al Sahili et al., 2024); and VHELM offers a multi-dimensional benchmark for bias, fairness, toxicity, and safety (Lee et al., 2024). Finally, Janghorbani & De Melo (2023) introduces a unified framework for stereotypical bias across modalities, and Raj et al. (2024) exposes hidden biased associations that elude existing diagnostics.
@@ 2.3 Scaling Effects on Bias
Scaling model parameters or data volume is often viewed as a route to robustness, yet empirical findings remain mixed. In language models, larger GPT-style systems sometimes reduce toxic generation (Hoffmann et al., 2022) but can entrench occupational stereotypes (Liang et al., 2021). Ghate et al. (2025) shows that intrinsic bias is largely predictable from pre-training data, not architecture. Liu et al. (2024) offers a probabilistic framework revealing that stereotype volatility increases with model size. For VLMs, Birhane et al. (2023) warned that expanding LAION from 400M to 5B images risks amplifying existing harms, whereas higher image resolution can mitigate certain gender biases. Our controlled study extends this line by showing that size, scale, and loss source interact non-linearly: the same parameter increase that softens bias in CLIP amplifies it in OpenCLIP, and a five-fold data increase leaves CLIP’s skew almost unchanged while doubling that of OpenCLIP.
In summary, prior work establishes that VLMs inherit social bias, offers diverse metrics and mitigation heuristics, and yields conflicting evidence on the benefits of scaling. We advance the field by isolating the individual and joint contributions of model size, data scale, and loss source—an experimental control that existing audits lack.
@@ 3 Methodology
Our aim is to isolate how three levers—encoder size, pre training scale, and training data source—shape social bias in contrastive vision–language models (VLMs).
All other ingredients, most notably the symmetric cross-entropy loss that underpins CLIP, are held constant so that any change in measured bias can be attributed to those three levers alone.
The design forms a fully crossed $2\times 2\times 2$ grid (ViT B/32 or ViT L/14; 400 M or 2 B image–text pairs; proprietary or LAION style corpus), yielding eight candidate checkpoints.
@@
Where a cell is missing, we mark it explicitly and confine statistical comparisons to the subset of cells that share all three settings.
3.1 Background: Contrastive Pre-training
Both CLIP and OpenCLIP train a Vision Transformer $f_{\theta}$ and a text Transformer $g_{\phi}$ to align images $I$ and captions $t$ in a shared embedding space. For a batch of $N$ paired samples $\{(I_{i},t_{i})\}_{i=1}^{N}$ the models minimise the symmetric cross entropy
@@ 3.2 Evaluation Datasets
FairFace contains 108 501 cropped, face-only portraits labelled for seven self-identified race categories {White, Black, Indian, East Asian, South East Asian, Middle Eastern, Latino} and binary gender, sampled from Flickr under a CC BY–NC licence (Karkkainen & Joo, 2021). We draw the validation subset of 10 954 portraits such that every race $\times$ gender combination contains $\approx$ 782 images. This face-only design provides no scene context.
PATA.
The Protected Attribute Tag Association (PATA) benchmark comprises 4 934 images of people organised into 24 scenes (e.g., office, lab, sports), each annotated with binary gender (male/female) and five ethno-racial identities {Black, Caucasian, East Asian, Hispanic/Latino, Indian(Seth et al., 2023). Unlike FairFace, PATA includes scene context and curated caption sets designed to probe positive/negative associations (78 unique positive and 88 negative captions in total), enabling context-sensitive bias evaluation.
3.3 Bias Probing Tasks
Crime vs. Non human.
Following the original CLIP paper (Radford et al., 2021), we contrast ‘‘photo of a <label> criminal’’ with ‘‘photo of a <label> animal’’, where <label> instantiates the protected attribute under study (e.g., “Black man”, “Asian woman”). The template covers gender, race, and a Non human catch-all to test whether models associate certain demographics more strongly with criminality.
Communion and Agency.
To probe higher-level social stereotypes, we adopt the adjective lists introduced by Hausladen et al. (2024). For each trait pair $(\textsc{Positive},\textsc{Negative})$ on the communion and agency axes, we build the caption ‘‘photo of a <trait> person’’. This setup isolates representational bias in the vision-language association itself; no ground truth labels are required.
3.4 Bias Metrics
Setup.
Let $a$ be a protected attribute (e.g., Gender, Race) with groups $\mathcal{G}_{a}$. For group $g\in\mathcal{G}_{a}$, $D_{g}$ is the set of images in that group. Given image $I$ and caption $c$, we compute cosine similarity
$s(I,c)=\cos\!\bigl{(}v(I),e(c)\bigr{)},$
where $v(\cdot)$ and $e(\cdot)$ are L2-normalised image and text embeddings. For each bias-probing task tt3.3), we define a set of events $\mathcal{E}_{t}$ (e.g., Criminal, negative Communion).
Outcome proportion.
The proportion of group $g$ whose top-1 prediction falls in event $E\in\mathcal{E}_{t}$ is
$p_{g}(E)=\frac{1}{|D_{g}|}\sum_{I\in D_{g}}\mathbb{1}\!\left[\hat{c}(I)\in E\right],$
where $\hat{c}(I)=\arg\max_{c\in\mathcal{C}_{t}}s(I,c)$ and $\mathcal{C}_{t}$ is the candidate set for task $t$.
Max Skew (disparity).
We compute Max Skew separately for Race and Gender, then report the mean Max Skew across all unordered group pairs for race (mean max skew for race) and the single pair for gender. For two groups $A,B\in\mathcal{G}_{a}$ and event $E\in\mathcal{E}_{t}$,
$S_{A,B}(E)=\max\left(\left|\frac{p_{A}(E)-p_{B}(E)}{p_{B}(E)}\right|,\;\left|\frac{p_{B}(E)-p_{A}(E)}{p_{A}(E)}\right|\right),$
captures the largest relative gap in selection rates between $A$ and $B$, regardless of which is higher. We then average over all events and relevant group pairs:
$\overline{S}(a,t)=\frac{1}{|\mathcal{E}_{t}|\binom{|\mathcal{G}_{a}|}{2}}\sum_{E\in\mathcal{E}_{t}}\sum_{\begin{subarray}{c}A,B\in\mathcal{G}_{a}\\ A<B\end{subarray}}S_{A,B}(E).$
For instance, in the Crime probing task, if 12% of portraits of Black men are tagged Criminal but only 8% of White men are, the relative differences are (0.120.08)/0.08=0.50(0.12-0.08)/0.08=0.50 (50%) and (0.080.12)/0.12=0.33(0.08-0.12)/0.12=-0.33 (33% in magnitude); Max Skew takes the larger magnitude, $0.50$, meaning the model assigns the Criminal label 50% more often to Black men than to White men for that task. Lower $\overline{S}(a,t)$ indicates less disparity, with 0 denoting perfect parity.
Harm Rate (incidence).
The Harm Rate for event $E$ is the overall fraction of images predicted with that label:
$h(E)=\frac{1}{|\mathcal{I}|}\sum_{I\in\mathcal{I}}\mathbb{1}\!\left[\hat{c}(I)\in E\right].$
This measures how often harmful predictions occur, regardless of which group they affect. We report $h(E)$ alongside $\overline{S}(a,t)$ for each task.
3.5 Debiasing Strategies
To probe whether the scale factors interact with explicit mitigation techniques, we apply three recently-proposed methods that operate on frozen CLIP encoders:
Bias Prompts (projection & calibration).
We adopt the training-free approach of Chuang et al. (2023), which constructs a linear projection that removes protected attribute directions from text features only. Concretely, we form a matrix $A$ whose columns are CLIP embeddings of attribute prompts (e.g., male/female, race terms) and compute the orthogonal projector $P_{0}=I-A(A^{\top}A)^{-1}A^{\top}$. To stabilise purely prompt-defined subspaces, we follow the authors’ calibration step: a closed-form objective encourages projected pairs such as “male doctor” and “female doctor” to coincide, yielding a calibrated matrix $P^{\star}$ that we then apply to all evaluation captions. This method requires no training data, labels, or changes to the image encoder.
Prompt Array (adversarial prompt tuning).
We implement the adversarial debiasing scheme of Berg et al. (2022a), prepending a small array of learnable tokens to sensitive text queries (our crime/communion/agency prompts). The tokens are trained so that an adversary fed only image–text similarity scores cannot predict the protected attribute; a simultaneous image–text contrastive term preserves the joint representation. This protocol uses attribute labels during training for the adversary but does not modify backbone weights.
SANER (annotation-free neutralizer).
@@ 3.5 Debiasing Strategies
Compatibility remark.
Only BP functions on both the closed-weight OpenAI CLIP and the open-weight OpenCLIP models; PA and SANER require back-propagation through the text encoder and are therefore only evaluated on the open-source models.
@@ 3.6 Inference Details
Images are resized to $224\times 224$ for ViT B/32 and $336\times 336$ for ViT L/14 to match pre-training. Caption embeddings use the checkpoint-specific temperature $\tau$ without test time augmentation. BP uses the authors’ original prompt pairs and the calibrated projection matrix released with the paper. PA is trained for three epochs on 90k FairFace images plus the same number of LAION images, using the hyperparameters from the AACL-22 code release ($\lambda_{\text{ITC}}{=}0.05$). SANER follows the SD-XL caption protocol of Hirota et al. (2025) and is trained for five epochs on COCO 2017. During inference, the debias layer or token array is inserted after all pre-processing so that the downstream evaluation pipeline (§LABEL:ssec:_tasks) remains identical. All three methods operate without retraining the backbone encoders. Bias Prompts and SANER adjust text features at inference (closed form or light layer), while Prompt Array learns a small set of tokens with an adversary to reduce attribute leakage. A single NVIDIA A100 (40 GB) processes the full benchmark in under thirty minutes. Code, prompts, and raw outputs will be released upon publication.
@@ 4 Results
This section traces how encoder size, pre-training scale, and training-data source alter social bias, and evaluates the extent to which post-hoc debiasing can attenuate these effects. We benchmark three test-time strategies—Bias Prompts, Prompt Array, and SANER—across crime, communion, and agency for gender and race. All baseline numbers come from Table 1; Figures 1, 2, and 3 visualise the same factors, and mitigation results are summarised in §4.6.
Data Model Sz DSz $\max s_{G}^{c}$ $\max s_{G}^{com}$ $\max s_{G}^{ag}$ $\mu\max s_{R}^{c}$ $\mu\max s_{R}^{com}$ $\mu\max s_{R}^{ag}$ % C % NH % NC % NA
@@
Table 2: Bias component deltas across model size, data size, and data source. Red = increase in bias (undesirable), Blue = reduction in bias (desirable).
Data Model Debias Sz DSz $\max s_{G}^{c}$ $\max s_{G}^{com}$ $\max s_{G}^{ag}$ $\mu\max s_{R}^{c}$ $\mu\max s_{R}^{com}$ $\mu\max s_{R}^{ag}$ % C % NH % NC % NA
Fair CLIP biased prompts B/32 400M 0.11 0.07 0.05 2.95 0.17 0.48 28.18 0.24 45.37 63.04
biased prompts L/14 400M 1.09 0.18 0.03 0.98 0.21 0.16 30.03 0.12 65.46 74.77
OCLIP saner B/32 400M 0.21 0.82 0.11 1.49 0.70 0.15 9.88 0.34 8.38 56.07
saner L/14 400M 2.84 0.72 0.43 1.85 0.34 0.18 3.88 0.37 24.46 63.30
biased prompts B/32 400M 0.47 0.23 0.01 1.75 0.55 0.05 14.33 1.90 20.37 82.76
prompts array B/32 400M 0.71 0.52 0.03 1.51 0.46 0.09 5.74 0.71 9.67 77.06
biased prompts L/14 400M 0.12 0.71 0.10 4.31 0.35 0.16 9.00 3.44 20.37 82.76
prompts array L/14 400M 2.23 0.84 0.34 2.14 0.37 0.19 3.34 0.47 16.46 77.06
biased prompts B/32 2B 0.59 0.29 0.01 1.95 0.29 0.04 17.09 0.09 41.84 81.80
prompts array B/32 2B 1.12 0.33 0.09 1.51 0.24 0.04 5.23 0.03 41.84 81.80
saner B/32 2B 1.89 0.27 0.05 0.90 0.19 0.11 7.66 0.08 37.25 74.93
biased prompts L/14 2B 0.79 0.55 0.04 6.84 0.46 0.10 13.58 0.63 12.46 43.19
prompts array L/14 2B 1.96 0.52 0.36 6.64 0.36 0.20 2.81 0.11 16.46 29.03
saner L/14 2B 1.92 0.54 0.45 2.93 0.39 0.25 3.28 0.47 18.35 84.64
PATA CLIP biased prompts B/32 400M 0.16 0.13 0.31 1.21 0.17 0.22 10.28 0.30 28.14 30.07
biased prompts L/14 400M 3.47 0.49 0.00 4.76 0.27 0.16 8.21 0.28 12.46 43.19
OCLIP biased prompts B/32 400M 0.75 0.29 0.06 0.94 0.35 0.14 15.45 2.84 15.27 34.41
prompts array B/32 400M 1.78 0.02 0.02 0.93 0.38 0.31 6.91 1.06 11.30 34.42
biased prompts L/14 400M 0.21 0.17 0.02 2.53 0.50 0.16 22.09 3.37 17.65 29.61
prompts array L/14 400M 2.28 0.00 0.22 2.21 0.47 0.13 6.69 0.73 16.46 29.03
saner B/32 400M 1.21 0.06 0.05 1.05 0.25 0.29 7.45 0.79 11.25 32.22
saner L/14 400M 3.70 0.04 0.25 1.84 0.43 0.13 6.41 0.86 18.24 32.22
biased prompts B/32 2B 0.07 0.34 0.32 2.43 0.55 0.13 19.35 0.26 15.27 34.41
prompts array B/32 2B 1.12 0.01 0.07 1.66 0.19 0.22 5.12 0.03 11.30 34.42
saner B/32 2B 1.50 0.04 0.01 1.41 0.33 0.19 6.00 0.03 12.41 32.04
saner L/14 2B 1.26 0.25 0.18 5.00 0.20 0.13 7.02 0.15 15.45 26.49
prompts array L/14 2B 1.07 0.31 0.13 5.33 0.25 0.11 6.16 0.15 14.11 22.90
biased prompts L/14 2B 0.16 0.34 0.11 3.48 0.27 0.09 23.02 0.68 17.38 26.11
Table 3: Bias and toxicity metrics for de-biased CLIP/OpenCLIP models. Blue marks the best score, red the worst within each metric.
@@
Table 4: Race-related bias deltas across model size, data size, and dataset source. Red = increase in bias; Blue = decrease in bias.